### v1

# LLM Reading Tea Leaves: Automatically Evaluating Topic Models with Large Language Models

Abstract
Topic modeling has been a widely used tool for unsupervised text analysis. However, comprehensive evaluations of a topic model remain challenging. Existing evaluation methods are either less comparable across different models (e.g., perplexity) or focus on only one specific aspect of a model (e.g., topic quality or document representation quality) at a time, which is insufficient to reflect the overall model performance. In this paper, we propose WALM (Words Agreement with Language Model), a new evaluation method for topic modeling that comprehensively considers the semantic quality of document representations and topics in a joint manner, leveraging the power of large language models (LLMs). With extensive experiments involving different types of topic models, WALM is shown to align with human judgment and can serve as a complementary evaluation method to the existing ones, bringing a new perspective to topic modeling. Our software package will be available at https://github.com/Xiaohao-Yang/Topic_Model_Evaluation, which can be integrated with many widely used topic models.
1 Introduction
Topic modeling Blei et al. (2003), a popular unsupervised text analysis technique, has been applied to various domains, including information retrieval (Yi and Allan, 2009), marketing analysis (Reisenbichler and Reutterer, 2019), social media analysis (Laureate et al., 2023), bioinformatics (Liu et al., 2016), and more. A topic model typically learns a set of global topics to interpret a text corpus and the topic proportion of a document as its semantic representation.
Although topic models have been time-tested for two decades, as an unsupervised technique, comprehensive evaluations of a topic model remain challenging (Zhao et al., 2021). Originally, topic models are implemented as probabilistic graphical models such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and many of its Bayesian extensions. For these models, it has been common practice to measure the log-likelihood or perplexity of a model on held-out test documents. While log-likelihood or perplexity provides a straightforward quantitative comparison between models, several issues still persist: As topic models are not primarily designed for predicting words in documents but rather for semantically learning interpretable topics and document representations, these metrics do not measure these qualities. Furthermore, estimating the predictive probability is often intractable for Bayesian models, and different papers may employ different sampling or approximation techniques (Wallach et al., 2009; Buntine, 2009). For recently proposed neural topic models (NTMs) (Zhao et al., 2021), the computation of log-likelihood is even more inconsistent.
In addition to log-likelihood or perplexity, document representation quality and topic quality are evaluated separately. For document representation quality, downstream task performance is typically used as a metric, such as document classification (Yang et al., 2023), clustering (Zhao et al., 2021), and retrieval (Larochelle and Lauly, 2012). For topic quality, the ultimate evaluation method is human evaluation, which is time-consuming and expensive. Thus, various automatic metrics have been proposed, such as topic coherence (Lau et al., 2014), which measures how semantically coherent the representative words in a topic are, and topic diversity (Dieng et al., 2020), which measures how diverse discovered topics are. To comprehensively evaluate the performance of a topic model, one needs to report multiple metrics on both document representation and topic qualities. However, these metrics can be contradictory; e.g., a topic model with good topic quality may not preserve good quality on document representation, and vice versa. This discrepancy complicates the model selection process for topic models in practice.
In this paper, we aim to develop a new evaluation approach for topic modeling that comprehensively considers the semantic quality of document representations and topics in a joint manner, leveraging the power of large language models (LLMs). Our key idea is as follows: After being trained, a topic model can infer a document’s distribution over topics and each topic is a distribution over vocabulary words. With these two distributions, a model can generate a set of “topical” words given a document, such as by looking at its representative topics and the representative words of each topic. The generation of the topical words takes both the topic distribution of a document and the word distributions of the topics into account, which captures the semantic summary of the document and is expected to align with the keywords identified by humans. Thus, the quality of the generated topical words of documents informatively and comprehensively reflects the quality of a topic model. As human evaluation is typically costly, we propose querying LLMs with appropriate prompts as a proxy to generate keywords for the document and then compare them with the topical words generated by a topic model. Finally, to quantify the agreement between the words from the topic model and the LLM, a series of WALM (Words Agreement with Language Model) metrics are proposed. WALM has the following appealing properties:
It is a joint metric that evaluates both document representation and topic qualities.
It measures how well a topic model captures a document’s semantics, which is the focus of topic modeling.
It can be compared across different kinds of topic models.
To examine WALM series metrics, we conduct extensive experiments using various popular topic models on different datasets, comparing them with other widely used topic model evaluation metrics. Moreover, human evaluation is also conducted to demonstrate the alignment of WALM with human judgment.
2 Related Work
As an unsupervised technique for uncovering hidden themes in text, evaluating topic models remains challenging. Early evaluations of a topic model rely on the log-likelihood or perplexity of held-out documents (Blei et al., 2003), which measures how well the model predicts the words of documents. As the computation of predictive probability is often intractable for conventional Bayesian topic models, various sampling or approximation techniques have been proposed Wallach et al. (2009); Buntine (2009). Apart from the inconsistent estimation, held-out likelihood is regarded as not correlated with the interpretability of topics from a human perspective (Chang et al., 2009), prompting the direct evaluation of topics and document representation quality.
As for the evaluation of topics, Chang et al. (2009) design the word and topic intrusion tasks for human annotators, where high-quality topics or document representations are those where annotators can easily identify the intruders. Newman et al. (2010) and Mimno et al. (2011) evaluate topic coherence by direct ratings from human experts. Although human judgment is commonly regarded as the gold standard, it is expensive and impractical for large-scale evaluation. Automated evaluation of topic coherence is more practical, such as Normalized Pointwise Mutual Information (NPMI) (Lau et al., 2014), which relies on the co-occurrence of the topic’s top words in the reference corpus to measure topic coherence, with the underlying assumption that a large reference corpus such as Wikipedia can capture prevalent language patterns. Although they automate the evaluation of topics and strongly correlate with human judgment (Newman et al., 2010), counting word co-occurrence in a large reference corpus is still relatively expensive. Recent works propose leveraging word embeddings (Nikolenko, 2016) or contextualized embeddings (Hoover et al., 2021) for efficiently evaluating topic coherence, incorporating semantics from pre-trained embeddings. Due to common posterior collapse issues (Lucas et al., 2019) in the growing field of neural topic models (Zhao et al., 2021), recent works also consider topic diversity (Dieng et al., 2020) during evaluation, which measures how distinct the top words of each topic are.
As for the evaluation of document representation, early works focus on how well the topic proportion of a document represents the document content, assessed through a topic intrusion task by human annotators (Chang et al., 2009), which is further extended as automated metrics (Bhatia et al., 2017, 2018). Recent topic models often use the topic proportions as document representations, the quality of which is commonly investigated through downstream tasks, including their use as features for document classification (Nguyen and Luu, 2021), clustering (Zhao et al., 2020), and retrieval (Larochelle and Lauly, 2012). Recently, the generalization ability of topic models is investigated by evaluating their quality of document representations across different unseen corpora (Yang et al., 2023).
In the era of large language models (LLMs) (Thoppilan et al., 2022; Chowdhery et al., 2022; OpenAI, 2022; Touvron et al., 2023a, b), recent research has begun leveraging LLMs to evaluate topic models, such as using ChatGPT (OpenAI, 2022) as a proxy for human annotators for word intrusion and topic rating tasks for evaluating topic coherence (Rahimi et al., 2023; Stammbach et al., 2023). The focus of these works is still on topic quality only.
In this work, we propose new evaluation metrics for topic models, differing from previous works in the following ways: (1) Unlike evaluations that focus on only sub-components of a topic model (i.e., topics or document representations), our evaluation metrics offer a joint approach to topic model evaluation, considering both topics and document representations together. (2) Compared with log-likelihood or perplexity, which also evaluate based on documents, our evaluation metrics consider semantics from documents and align with the focus of topic modeling. (3) Different from recent LLM-based evaluations that use LLMs for topic quality evaluation, ours is a more comprehensive method that considers both topic quality and document representation quality and our use of LLMs is quite different from previous works.
3 Background
Given a document collection with vocabulary words, a topic model is typically trained on their Bag-of-Words (BOWs), e.g., . The topic model can infer a distribution over topics for each document by running its inference process:
where denotes the model parameters of the inference process; ( denotes the probability simplex) indicates the proportion of each topic present in the document and is commonly used as its semantic representation. Besides, the topic model also discovers global topics for the corpus (i.e., ), where each topic is a distribution over vocabularies. Ideally, each topic captures a semantic concept that can be interpreted by its top-weighted words. To train a topic model, one often needs to generate or reconstruct the word distribution of the document from by running its generative process:
where are the model parameters of the generative process; is the per-document word distribution from which is sampled. Let be the semantic representations of test documents and be the learned topics, current evaluation of a topic model is commonly conducted based on either or separately.
4 Method
4.1 Motivation
Both topics and document representations are important components of a topic model. To comprehensively evaluate a topic model, it is common practice to report the performance of both parts. This can be done by measuring topic quality using metrics such as NPMI and assessing document representation quality through downstream classification accuracy (ACC) (see section 5.1 for details of metrics calculation). However, a model that prioritizes topic quality (e.g., NPMI) may not perform well in terms of document representations (e.g., ACC), and vice versa, which creates difficulty during model selection, as illustrated in Figure 1. This inconsistency in the performance of the two components is also indicated by Bhatia et al. (2017). Therefore, evaluating a topic model based on sub-components only is insufficient to reveal the entire model’s performance. Recent topic models usually focus on improving topic quality, such as clustering-based topic models in Sia et al. (2020); Grootendorst (2022), but they do not evaluate their performance for document representation. In this work, we aim to introduce a novel evaluation method for topic modeling that jointly assesses the semantic quality of both topics and document representations, with the help of large language models.
4.2 Key Idea
We propose to conduct the evaluation in a joint manner that considers both document representations and topics, rather than evaluating them separately as in previous works. To do so, we obtain the document-word distribution for a given document from the topic model by running both its inference and generative process:
The inference process infers the document representation for a given document , as in Eq. 1; the generative process generates or reconstructs the word distribution based on the document representation and topics , as in Eq. 2. Therefore, the evaluation based on involves both document representations and topics.
Next, we take the top-weighted words from the word distribution generated by the topic model as the “topical” words of the document. Those topical words can be regarded as a semantic summary of the document from the target topic model’s perspective. To generate high-quality topical words for a document, a topic model should learn good global topics as well as good document representations. Now, the evaluation of the topic model can be converted to the evaluation of the quality of the topical words. Suppose the true representative words of document are given, then we can formulate our evaluation task as:
where is a score function (Sec. 4.4) to quantify the agreement between and (Sec. 4.3).
4.3 Document summarization by LLM
Following our evaluation task in Eq. 4, the ideal representative words are from human summary of the document. However, this is expensive and impractical for large-scale evaluation. With the recent advancements in LLMs, which have demonstrated performance akin to human capabilities in various NLP tasks, including text summarization Wang et al. (2023); Tang et al. (2023); Zhang et al. (2024), we propose to leverage LLMs by prompting to extract key words of documents:
Specifically, we query keywords for a given document from an LLM by proposing the prompt shown in Figure 2. The prompt consists of a task instruction, an in-context example (Liu et al., 2021), and the queried document.
4.4 Choices of the Score Function
For the score function in Eq. 4, we propose different ways to calculate it: (1) Overlap-based, which computes the number of overlapping words between and , and (2) Embedding-based, which calculates the overall semantic similarity between the two word sets using pre-trained word embeddings.
Word Overlap
A straightforward choice of the score function is directly counting the overlaps between and . Considering the potential variant in forms of the same word, we convert each word to its root form before counting, formulated as:
where and are the counting and rooting (e.g., stemming or lemmatization) operation, respectively; returns the normalising factor based on two input word sets, where and are the number of words in and , respectively.
Synset Overlap
Considering the case that different words may describe the same or similar concept (e.g., both “puppy” and “dog”), we leverage the synset of WordNet (Miller, 1995) when counting the number of overlaps: if there is an overlap between two words’ synsets, then it is considered as an overlap, formulated as:
where denotes the indicator function; is a function that returns the synset for a given word.
Word Optimal Assignment
We consider another choice of , which measures the overall semantic similarity between two sets of words with pre-trained word embeddings (Mikolov et al., 2013; Pennington et al., 2014). Since the alignments between words from and are unknown, directly measuring similarity between word embeddings is not feasible. To automatically find the alignment for each word of to each word of , we formulate it as the following optimal assignment (OA) problem and solve it using the Hungarian algorithm Kuhn (1955): Given a word set that has words: and their embedding vectors ; and another word set that has words: with related embedding vectors . Define a cost matrix whose entry , where denotes the cosine distance function; and a binary matrix whose entry if word is assigned to word , and 0 otherwise. The goal is to solve the following optimal assignment problem:
| subject to: | ||||
By finding the optimal binary matrix , we obtain the distance between and by:
Word Optimal Transport
Considering that both and are top words of probability distributions, where each word essentially retains a portion of probability mass. Our previous calculations ignore the probability mass of words and treat each word in the set as equal. Now we include the probability mass and formulate the similarity calculation between and as an optimal transport (OT) problem: Given two discrete distributions and , where and are the supports of those two distributions; and are their related probability vectors222We assume is a uniform distribution over the keywords from the LLM. Thus, is a uniform probability vector.; following the same construction of cost matrix in the previous OA problem, the OT problem between and is defined as:
| subject to: | ||||
where is the transport plan, whose entry indicates the amount of probability mass moving from to . Similarly, by finding the optimal transport plan using solvers such as those in Flamary et al. (2021), the OT distance between and is obtained by:
Compared with our OA and OT formulations for WALM, they are similar in that they both construct the cost matrix using cosine distance between pre-trained word embeddings. However, they differ in the following ways: (1) OA treats words in the set as equal, while OT considers probability mass of each word. (2) OA can be viewed as a “hard” assignment problem between two word sets because the entries of are binary. In contrast, OT can be regarded as a “soft” assignment because of the spread of probability mass in .
5 Experiments
5.1 Experimental Setup
Datasets
Evaluated Models
We conduct experiments on 7 popular topic models from traditional probabilistic to recent neural topic models. (1) Latent Dirichlet Allocation (LDA) (Blei et al., 2003), the most popular probabilistic topic model that assumes a document is generated by a mixture of topics. (2) LDA with Products of Experts (PLDA) (Srivastava and Sutton, 2017), an early NTM that applies the product of experts instead of the mixture of multinomials in LDA. (3) Neural Variational Document Model (NVDM) (Miao et al., 2017), a pioneer NTM that uses a Gaussian as the prior distribution of topic proportions of documents. (4) Embedded Topic Model (ETM) (Dieng et al., 2020), an NTM that involves word and topic embeddings in the generative process. (5) Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2017), an NTM that applies a logistic normal prior on topic proportions and leverages extra information from metadata. (6) Neural Sinkhorn Topic Model (NSTM) (Zhao et al., 2020), a recent NTM based on an optimal transport framework. (7) Contrastive Learning Neural Topic Model (CLNTM) (Nguyen and Luu, 2021), a recent NTM that uses contrastive learning to regularize document representations. We keep all these models’ default settings as suggested in their implementations. All experiments are conducted 5 times with different random seeds; mean and std values are reported.
Settings of WALM
We evaluate topic models using WALM series metrics we propose in Sec. 4.4, including , , , and . As for the settings of WALM, we use the word embeddings from GloVe (Pennington et al., 2014) pre-trained on Wikipedia333https://nlp.stanford.edu/projects/glove/. We set the number of top words for a document generated by the topic model to 10; The LLM we use is LLAMA2-13b-chat (Touvron et al., 2023b), which has 13 billion parameters and is fine-tuned for chat completion; We use a 4-bit quantization444https://huggingface.co/TheBloke/Llama-2-13B-chat-GGML version of this model for memory efficiency and leverage the llama.cpp555https://github.com/abetlen/llama-cpp-python library for easy inference. Notably, our evaluation metrics can be used with any LLMs. We use LLAMA2 in our study because it is open-sourced and cost-efficient.
Settings of Existing Metrics
We also evaluate topic models with existing commonly used metrics to compare with ours. (1) Topic Coherence and Diversity: we apply NPMI to evaluate topic coherence using Wikipedia as the reference corpus, done by the Palmetto package666https://github.com/dice-group/Palmetto (Röder et al., 2015). Following standard protocol, we consider the top 10 words of each topic and obtain the average NPMI score of topics by selecting the top 50% coherent topics. As for topic diversity (TD), we compute the percentage of unique words in the top 25 words of all topics, as defined in Dieng et al. (2020). (2) Document Representation Quality: we conduct document classification and clustering to evaluate the representation capability of topic models. As for classification, we train a Random Forest classifier based on the training documents’ representation and test the accuracy (ACC) in the testing documents, as in previous works such as Nguyen and Luu (2021). As for clustering, we conduct K-Means clustering based on test documents’ representation and report the Purity (KM-Purity) and Normalized Mutual Information (KM-NMI), as in previous works such as Zhao et al. (2020). (3) Perplexity: we use document completion perplexity (Wallach et al., 2009) to evaluate the predictive ability of topic models. We split each test document into two equal-length folds randomly. Then we compute the document completion perplexity (DC-PPL) on the second fold of documents based on the topic proportion inferred from the first fold, as in previous works such as Dieng et al. (2020).
| Document | Model | Topical Words |
[…]
| LDA_C | drive disk scsi hard card controller mb floppy ide sale | |
| NVDM_B | driver, drive, problem, card, time, file, thanks, need, email, work | |
| NVDM_C | drive, driver, hard, scsi, window, cd, mb, floppy, disc, work | |
| LLAMA2 | driver format magneto-optical disc operating system communication incompatibility | |
| Human | driver disc drive computer hardware software memory formatting format incompatible | |
[…]
| LDA_C | film played directed baseball league australian major drama football award | |
| NVDM_B | specie, album, school, known, located, north, film, directed, american, released | |
| NVDM_C | film, album, released, second, south, new, directed, american, australian, known | |
| LLAMA2 | wrong world film australian directed ian pringle victoria melbourne | |
| Human | film movie directed directer australian melbourne victoria |
5.2 Results and Analysis
Topic Model Evaluation with WALM
We assess topic models’ performance based on our evaluation metrics on both 20News and DBpedia. We have the following observations based on our results illustrated in Figure 3: (1) The WALM values of most models on DBpedia show better performance than 20News, which indicates that it is easier for topic models to generate informative topical words for short documents than long documents. (2) The performance indicated by overlap-based metrics (e.g., and ) and embedding-based metrics (e.g., and ) is slightly different. The reason is that embedding-based metrics consider the semantic distance among words, which can be more flexible than the exact match in overlap-based metrics. (3) It can be observed that there is little improvement from recent NTMs over LDA and NVDM in terms of our joint metrics. The potential reason is that most contemporary NTMs primarily focus on enhancing topic coherence while neglecting the generation of documents, thus showing weak performance in generating topical words of documents as indicated by WALM. (4) WALM shows consistent performance ranking across different settings of (e.g., and ) for long documents, which potentially alleviates the effort of extensive model selection in terms of during model comparison.
Learning Curves of WALM
We illustrate the learning curves of topic models in terms of WALM in Figure 4, which clearly shows how our metrics change along with the learning of a topic model. It demonstrates that our metrics are informative for model selection.
Qualitative Analysis on Topical Words for Documents
We qualitatively investigate the topical words of documents by topic models at different stages in Table 1, where we randomly sample one document for 20News and DBpedia, respectively. We have the following observations based on our results: (1) The topical words at the beginning phase contain less semantically related words about the documents than those at convergence, which aligns with the learning status (as in Figure 4) indicated by WALM. (2) The topical words of NVDM include more words that reveal the documents’ main messages than LDA, which aligns with the ranking (as in Figure 3) suggested by WALM. (3) The words from LLAMA2 are comparable with those from human annotators in the longer document. For the short document, LLAMA2 focuses on detailed information such as the movie title (e.g., “wrong” and “world”), which human annotators ignore. Further analysis between using LLAMA2 and human judgment to obtain the “true” topical words is described in Sec. 5.4.
Correlation to Other Metrics
We compute Pearson’s correlation coefficients among existing and WALM series metrics, similar to the correlation analysis in previous works such as Doogan and Buntine (2021) and Rahimi et al. (2023). Pearson’s correlation coefficients among the metrics are plotted in a heatmap in Figure 5. Based on the results, we observe that: (1) WALM variants are highly correlated with each other since they originate from the same mechanism. (2) Compared with DC-PPL, which also evaluates the entire model based on documents, WALM show weak correlations, suggesting a new family of evaluation metrics. (3) Compared with other types of evaluations, WALM has moderate correlations with document representation metrics (e.g., KM-Purity, KM-NMI, and ACC), and weak correlations with topic quality metrics (e.g., NPMI and TD). This indicates that our joint evaluation metrics consider both components. Those observations demonstrate the comprehensive nature of WALM, which can be used as a complementary evaluation method to the existing ones.
5.3 Contextualized Embeddings for WALM
Obtaining Contextualized Embeddings
Recall that in Eq. 9 and Eq. 4.4, the cost matrix is constructed using cosine distances between word embeddings. Here, we change our construction of from using static word embeddings from GloVe (Pennington et al., 2014) to the contextualized word embeddings from LLAMA2, considering that the same word may have different semantic meanings in different contexts. We obtain the contextualized embeddings of a word given a document differently in two cases: (1) When the target word appears in the context document, we take the average embeddings of each occurrence as the contextualized embedding. (2) When there is no occurrence of the target word in the given document, we add an auxiliary sentence to the document in the following format:
“<Given Document>. This document is talking about <Target Word>.”
Then, we obtain the contextualized embedding of the target word given the document with the auxiliary sentence. By replacing the global word embeddings with contextualized word embeddings, we have new variants of our embedding-based WALM (i.e., and ), i.e., and .
Observations
Since the cost of obtaining contextualized embeddings is high for LLMs, we compute and in a case study, where we test NVDM on 100 documents randomly sampled from the test sets of 20News and DBpedia, respectively. We plot the learning curves on test documents in Figure 6. We observe that using word embeddings or contextualized embeddings in our embedding-based scores exhibits similar trends but with different values on both datasets. Further analysis of both approaches is conducted in Sec. 5.4.
5.4 Comparisons with Human Evaluation
WALM computes the difference between documents’ topical words generated by topic models and an LLM, treating the words from the LLM as the ground truth. Here, we investigate the gap between using LLM and human judgment as the true topical words in WALM. To quantify this gap, we propose the following calculation:
where is the score function we propose in Sec. 4.4 (and the contextualized variants in Sec. 5.3); , , and are the topical words from the topic model, LLM and human annotators, respectively. We empirically observe that the topical words generated by topic models always differ from those identified by humans, thus the denominator in Eq. 12 will not be zero.
Case Study Setup
As human annotation is expensive for large-scale investigation, we randomly sample 200 test documents from 20News and DBpedia as a case study. We ask human annotators to provide keywords that capture the main points of each document. Then, given a trained topic model, we compute the gap between using the words from the LLM and human judgment in our metrics using Eq. 12.
Observations
The results are illustrated in Figure 7, where the evaluated model is NVDM with trained on 20News and DBpedia, respectively. We have the following observations based on the results: (1) Comparing the datasets, the gap between using human judgment and the LLM in 20News is lower than in DBpedia in most cases. This indicates that for long documents such as those in 20News, the topical words generated by the LLM are closer to human judgment than in short documents in DBpedia. (2) Comparing the metrics, exhibits the lowest gap among WALM metrics, with a gap value of 0.03 and 0.15 on 20News and DBpedia, respectively. This shows the effectiveness of using the LLM as a proxy for human judgment when applied in . (3) Comparing the embeddings, using contextualized embeddings from the LLM can further narrow the evaluation gap for and on short documents.
6 Conclusion
In this work, we propose WALM (Words Agreement with Language Model), a new evaluation method for topic models, which takes both topic and document representation quality into account jointly. WALM measures the agreement between the topical words generated by topic models and those from the LLM for given documents. We propose different calculations to quantify their agreement, including overlap-based and embedding-based metrics. Through extensive experiments, WALM series metrics show their effectiveness in reflecting topic models’ capability in terms of semantic summary of documents. We show that WALM metrics align with human judgment and can serve as an informative complementary method for topic model evaluation. In the future, we believe that WALM metrics can be further enhanced through improved document summarization by LLMs and word set similarity calculations, and topic models can be further optimized based on WALM series targets.